OpenStudy (anonymous):

A box is to be formed by cutting square pieces out of the corner of a rectangular piece of a 4" by 6" note card. The sides will then be folded up to make the box. A. the resulting length of the note card is 6-2x and the resulting width of the note card is 4-2. A(x)= (6-2x)(4-2x), set this equal to 16 and use the quadratic formula to solve for the two possible answers.

OpenStudy (anonymous):

A(x)= (6-2x)(4-2x) setting it 16 means 16 =(6-2x)(4-2x) 16 = (2*(3-x) ) * (2(2-x)) 16=4(3-x)(2-x) 4=(3-x)(2-x) and so on...

OpenStudy (anonymous):

i only got one answer when i solved it, so i'm not sure if i did something wrong.

OpenStudy (anonymous):

i got 5 +/- sqrt 17 over 2